# Augmenting Moment Retrieval: Zero-Dependency Two-Stage Learning

### v1

@@ Abstract
Existing Moment Retrieval methods face three critical bottlenecks: (1) data scarcity forces models into shallow keyword-feature associations; (2) boundary ambiguity in transition regions between adjacent events; (3) insufficient discrimination of fine-grained semantics (e.g., distinguishing “kicking” vs. “throwing” a ball). In this paper, we propose a zero-external-dependency Augmented Moment Retrieval framework, AMR, designed to overcome local optima caused by insufficient data annotations and the lack of robust boundary and semantic discrimination capabilities. AMR is built upon two key insights: (1) it resolves ambiguous boundary information and semantic confusion in existing annotations without additional data (avoiding costly manual labeling), and (2) it preserves boundary and semantic discriminative capabilities enhanced by training while generalizing to real-world scenarios, significantly improving performance. Furthermore, we propose a two-stage training framework with cold-start and distillation adaptation. The cold-start stage employs curriculum learning on augmented data to build foundational boundary/semantic awareness. The distillation stage introduces dual query sets: Original Queries maintain DETR-based localization using frozen Base Queries from the cold-start model, while Active Queries dynamically adapt to real-data distributions. A cross-stage distillation loss enforces consistency between Original and Base Queries, preventing knowledge forgetting while enabling real-world generalization. Experiments on multiple benchmarks show that AMR achieves improved performance over prior state-of-the-art approaches. Code is available at https://github.com/nagara214/AMR.
@@ 1 Introduction
The explosive growth of video platforms has spurred an urgent demand for the structured parsing of video content [68, 31, 38]. As one of the core tasks in video semantic understanding [3, 51, 74, 42], moment retrieval [40, 41, 28, 58] aims to locate precisely the start and end times of target segments in long video streams based on textual descriptions. Although this task demonstrates significant value in intelligent search [26] and video summarization [66, 78, 52], its development has been consistently constrained by the scarcity of annotated data and annotation ambiguity. Dense annotation of a single video segment requires substantial manual effort [2, 65, 12, 64], while the blurred boundaries between adjacent events (such as “preparing to shoot” and “completing the shot”) further exacerbate annotation inconsistency. Current mainstream methods often rely on additional data and pre-trained models to alleviate the data bottleneck, but these strategies typically necessitate the introduction of external knowledge or complex computational frameworks, leading to an imbalance between model generalization capability and resource overhead.
@@
Existing research primarily advances along three paths: feature alignment-based architecture optimization methods [71, 58] enhance local feature alignment by designing multimodal interaction modules, yet their constructed attention mechanisms exhibit limited discriminative power for semantically similar segments; pre-training-based knowledge transfer methods (e.g., SG-DETR [16]) leverage large-scale vision-language models to improve semantic representation, while external large-model-based data synthesis method [44] generate virtual video-text pairs, but the billion-level parameter counts result in skyrocketing deployment costs. These three approaches have failed to address a core contradiction: how to mine discriminative spatiotemporal reasoning patterns from limited annotated data under zero external dependencies?
@@
However, due to the inherent limitations of current single-stage training frameworks in effectively leveraging the combined benefits of augmented and real data, these models tend to converge to trivial solutions that merely capture mutation boundary characteristics from the augmented data. As shown in Figure 1, current single-stage training model (blue) exhibits significant performance limitations when applied to real-world data. To address this and achieve effective transfer of augmented capability, we construct a two-stage progressive training pipeline with a query-based transformer architecture [63, 59, 4, 41, 82, 6, 22, 23, 53, 54, 56, 60, 81]. The cold-start stage adopts a curriculum learning strategy, using augmented data to establish fundamental boundary and semantic discrimination capabilities in the model; The distillation adaptation stage innovatively designs a dual-path distillation mechanism with two distinct query sets, each with clearly defined roles: the original query set is designed to maintain boundary localization capability within the DETR architecture, while the active query set is designed to dynamically absorb real data characteristics. To prevent knowledge forgetting, cross-stage distillation loss functions are employed to constrain the distribution consistency between the original query and the base query, where the base query is trained from the cold-start model. This approach ensures a seamless transition and knowledge retention across stages, enabling the model to effectively adapt to real-world data while preserving its foundational capabilities learned from the cold-start stage (Figure 1 red).
@@ 2 Related Work
Moment Retrieval (MR) is a multi-modal grounding task [10, 33, 55, 61, 72, 19], aiming to localize video segments corresponding to textual queries. Early approaches fall into two categories: proposal-based methods [13, 1, 70, 9, 7, 79, 35], which rely on pre-generated candidate segments, and proposal-free methods [8, 36, 75, 14, 76, 48, 49], which directly predict temporal boundaries without proposals. Recent advancements leverage DETR-based architectures [29, 41, 59, 24, 27, 69, 63, 25, 62, 77], inspired by their success in object detection, to improve MR performance through end-to-end query modeling. Concurrently, multimodal large language models (MLLMs) [39, 37, 20, 67, 21, 18] exploit cross-modal knowledge to achieve remarkable results by aligning linguistic and visual semantics.
Beyond architectural innovations, data-centric approaches have emerged to address MR challenges. Some works [36, 8, 75, 50] tackle data imbalance issues or expensive computational costs, while others incorporate large-scale external datasets for pretraining. Notably, VIDEOLIGHTS [44] leverages BLIP-2’s image-to-text generation to synthesize training data. Although effective, these methods often depend on external resources. In contrast, we propose a simpler, generalizable data augmentation strategy that enhances performance without external data or pretrained models.
@@
Video Data Augmentation has been widely explored to address the challenges of limited and difficult-to-annotate video data. Common augmentation techniques involve spatial transformations, such as random flipping and cropping, or temporal modifications, such as random downsampling. In recent years, novel augmentation strategies have emerged. For instance, [73] introduced a method for video classification that randomly mixes videos along spatial and temporal dimensions, assigning ground truth labels proportionally based on voxel distribution. Building on this, [80] employed GAN [15] to generate more realistic video data, while [17] proposed a learning-based approach to select optimal video pairs for higher-quality augmentation. Additionally, [43] utilized video diffusion models to generate labeled videos for the guidewire segmentation task.
@@ 3 Method
Existing moment retrieval frameworks (e.g., M-DETR [29]) face significant challenges under data scarcity and distribution imbalance. While DETR-based architectures [59, 71, 69, 27] achieve end-to-end localization through encoder-decoder paradigms, their performance is constrained by two critical limitations: 1) High annotation costs for video data result in limited training samples, causing models to overfit specific training distributions; 2) Practical scenarios contain inherently ambiguous event boundaries with adjacent segment overlaps and semantically similar events that induce localization confusion.
@@
Our work focuses on leveraging limited and imbalanced training samples to enhance temporal localization models’ generalization capabilities in real-world scenarios. Consequently, the proposed AMR adheres to the standard vanilla DETR-based framework, comprising feature extraction, transformer encoder, and transformer decoder modules. For feature extraction, we align with previous work by using CLIP [46] and SlowFast [11] to extract video features while employing the CLIP text encoder for text feature extraction, where the encoded video features are denoted as $V\in\mathbb{R}^{L\times C}$ ($L$ video frames) and the text features as $T\in\mathbb{R}^{K\times C}$ ($K$ text tokens), with $C$ representing the hidden dimension. In the transformer encoder, cross-attention from visual to text features encodes multimodal video representation $\hat{V}\in\mathbb{R}^{L\times C}$ and text representation $\hat{T}\in\mathbb{R}^{K\times C}$. The decoder processes $N$ learnable queries $Q\in\mathbb{R}^{N\times C}$ with video features through several attention blocks and feed-forward networks, generating updated queries $\hat{Q}\in\mathbb{R}^{N\times C}$. Final boundary coordinates and confidence scores are predicted through bounding box and classification heads.
@@ 3.2 Two-stage Training: Cold Start and Distillation
In traditional single-stage training frameworks [59, 71, 69, 27], the distribution discrepancy between augmented data $\mathcal{D}_{\text{aug}}$ and real data $\mathcal{D}_{\text{real}}$ often leads to a decline in model generalization. Specifically, the sharp boundaries of events in synthetic data may cause the model to over-rely on abrupt visual features (e.g., scene cuts or sudden object appearances) while neglecting gradual transition patterns in real-world scenarios (e.g., natural continuity of human actions). To address this issue, we propose a cold-start distillation two-stage training framework (Fig. 3), whose core idea is to establish fundamental boundary localization and semantic modeling capabilities through augmented data in the cold-start stage and achieve migration to real data via a knowledge distillation strategy in the distillation adaptation stage.
@@ 3.3 Training Objective
Hungarian Matching To establish one-to-one correspondence between predicted and ground-truth objects without manual heuristics, we employ the Hungarian matching algorithm following the DETR framework [4]. The matching task is formulated as a minimum-cost bipartite graph matching problem. Specifically, a bipartite graph is constructed between the prediction set (containing $N$ elements) and the padded ground-truth set (expanded to $N$ elements with empty sets). The optimal assignment is determined by minimizing the matching cost:
@@
Loss Function During the cold-start stage training, we utilize the baseline loss functions from M-DETR [29], including the classification loss $\mathcal{L}_{\text{cls}}$, the localization loss $\mathcal{L}_{\text{loc}}$, and the auxiliary saliency loss $\mathcal{L}_{\text{sal}}$ (see [29] for implementation details). The composite loss function is defined as:
@@
M-DETR [29] 53.94 34.84 - - 32.20 52.89 33.02 54.82 29.40 30.73
UMT [34] 60.26 44.26 - - 38.59 56.23 41.18 53.83 37.01 36.12
QD-DETR [41] 62.68 46.66 62.23 41.82 41.22 62.40 44.98 62.52 39.88 39.86
UniVTG [32] 59.74 - - - 36.13 58.86 40.86 57.60 35.59 35.47
EaTR [24] 61.36 45.79 61.86 41.91 41.74 - - - - -
MomentDiff [30] - - - - - 57.42 39.66 54.02 35.73 35.95
TR-DETR [59] 67.10 51.48 66.27 46.42 45.09 64.66 48.96 63.98 43.73 42.62
TaskWeave [71] 64.26 50.06 65.39 46.47 45.38 - - - - -
UVCOM [69] 65.10 51.81 - - 45.79 63.55 47.47 63.37 42.67 43.18
BAM-DETR [27] 65.10 51.61 65.41 48.56 47.61 62.71 48.64 64.57 46.33 45.36
@@ 4.1 Datasets and Metrics
Datasets Our study evaluates performance on three benchmark datasets for moment retrieval: QVHighlights [29], Charades-STA [13], and TACoS [47]. QVHighlights consists of 10,148 videos, primarily sourced from YouTube, including vlogs and news clips. Notably, this dataset allows a single textual query to correspond to multiple segments within a video, making it more representative of real-world applications. Therefore, we adopt it as our primary benchmark. Charades-STA comprises 9,848 videos, focusing on human activities in indoor scenes. TACoS consists of 127 cooking-related videos, characterized by long video durations and short target segments, posing a significant challenge for moment retrieval.
@@
2D-TAN [79] 58.76 46.02 27.50 41.25 40.01 27.99 12.92 27.22
M-DETR [29] 65.83 52.07 30.59 45.54 37.97 24.67 11.97 25.49
MomentDiff [30] - 55.57 32.42 - 44.78 33.68 - -
QD-DETR [41] - 57.31 32.55 - - - - -
UniVTG [32] 70.81 58.01 35.65 50.10 51.44 34.97 17.35 33.60
UVCOM [69] - 59.25 36.64 - - 36.39 23.32 -
@@
Metrics We employ evaluation metrics consistent with prior works [71, 28]. For QVHighlights, we report Recall@1 (R1) at Intersection over Union (IoU) thresholds of 0.5 and 0.7. Given that this dataset allows multiple matching segments per query, we also use mean average precision (mAP) as an evaluation metric, reporting scores at IoU thresholds of 0.5, 0.75, and the average over all IoUs. For Charades-STA and TACoS, we evaluate Recall@1 (R1) at IoU thresholds of 0.3, 0.5, and 0.7, along with the mean IoU (mIoU) of the top-1 predictions.

### v2

@@ Abstract
Existing Moment Retrieval methods face three critical bottlenecks: (1) data scarcity forces models into shallow keyword-feature associations; (2) boundary ambiguity in transition regions between adjacent events; (3) insufficient discrimination of fine-grained semantics (e.g., distinguishing “kicking” vs. “throwing” a ball). In this paper, we propose a zero-external-dependency Augmented Moment Retrieval framework, AMR, designed to overcome local optima caused by insufficient data annotations and the lack of robust boundary and semantic discrimination capabilities. AMR is built upon two key insights: (1) it resolves ambiguous boundary information and semantic confusion in existing annotations without additional data (avoiding costly manual labeling), and (2) it preserves boundary and semantic discriminative capabilities enhanced by training while generalizing to real-world scenarios, significantly improving performance. Furthermore, we propose a two-stage training framework with cold-start and distillation adaptation. The cold-start stage employs curriculum learning on augmented data to build foundational boundary/semantic awareness. The distillation stage introduces dual query sets: Original Queries maintain DETR-based localization using frozen Base Queries from the cold-start model, while Active Queries dynamically adapt to real-data distributions. A cross-stage distillation loss enforces consistency between Original and Base Queries, preventing knowledge forgetting while enabling real-world generalization. Experiments on multiple benchmarks show that AMR achieves improved performance over prior state-of-the-art approaches. Code is available at https://github.com/SooLab/AMR.
@@ 1 Introduction
The explosive growth of video platforms has spurred an urgent demand for the structured parsing of video content [wu2016harnessing, lin2019tsm, lu2016optasia]. As one of the core tasks in video semantic understanding [buch2022revisiting, sadhu2021visual, zeng2020dense, nan2021interventional], moment retrieval [moon2023cgdetr, moon2023qd-detr, lei2021detecting, sun2024tr] aims to locate precisely the start and end times of target segments in long video streams based on textual descriptions. Although this task demonstrates significant value in intelligent search [korf1999artificial] and video summarization [wactlar2000informedia, zhang2016context, sharghi2016query], its development has been consistently constrained by the scarcity of annotated data and annotation ambiguity. Dense annotation of a single video segment requires substantial manual effort [bianco2015interactive, vondrick2013efficiently, galasso2013unified, voigtlaender2021reducing], while the blurred boundaries between adjacent events (such as “preparing to shoot” and “completing the shot”) further exacerbate annotation inconsistency. Current mainstream methods often rely on additional data and pre-trained models to alleviate the data bottleneck, but these strategies typically necessitate the introduction of external knowledge or complex computational frameworks, leading to an imbalance between model generalization capability and resource overhead.
@@
Existing research primarily advances along three paths: feature alignment-based architecture optimization methods [yang2024taskweave, sun2024tr] enhance local feature alignment by designing multimodal interaction modules, yet their constructed attention mechanisms exhibit limited discriminative power for semantically similar segments; pre-training-based knowledge transfer methods (e.g., SG-DETR [gordeev2024saliency]) leverage large-scale vision-language models to improve semantic representation, while external large-model-based data synthesis method [paul2024videolights] generate virtual video-text pairs, but the billion-level parameter counts result in skyrocketing deployment costs. These three approaches have failed to address a core contradiction: how to mine discriminative spatiotemporal reasoning patterns from limited annotated data under zero external dependencies?
@@
However, due to the inherent limitations of current single-stage training frameworks in effectively leveraging the combined benefits of augmented and real data, these models tend to converge to trivial solutions that merely capture mutation boundary characteristics from the augmented data. As shown in Figure 1, current single-stage training model (blue) exhibits significant performance limitations when applied to real-world data. To address this and achieve effective transfer of augmented capability, we construct a two-stage progressive training pipeline with a query-based transformer architecture [um2025kwdetr, sun2024trdetr, carion2020detr, moon2023qd-detr, zhu2025rethinking, chen2024survey12, huang2023free38, huang2025mvtokenflow39, shi2023edadet62, shi2023logoprompt63, shi2024plain66, tang2023contrastive70, zheng2023ddcot88]. The cold-start stage adopts a curriculum learning strategy, using augmented data to establish fundamental boundary and semantic discrimination capabilities in the model; The distillation adaptation stage innovatively designs a dual-path distillation mechanism with two distinct query sets, each with clearly defined roles: the original query set is designed to maintain boundary localization capability within the DETR architecture, while the active query set is designed to dynamically absorb real data characteristics. To prevent knowledge forgetting, cross-stage distillation loss functions are employed to constrain the distribution consistency between the original query and the base query, where the base query is trained from the cold-start model. This approach ensures a seamless transition and knowledge retention across stages, enabling the model to effectively adapt to real-world data while preserving its foundational capabilities learned from the cold-start stage (Figure 1 red).
@@ 2 Related Work
Moment Retrieval (MR) is a multi-modal grounding task [dai2024curriculum20, lin2021structured46, shi2024part2object64, tang2023temporal71, yang2021bottom81, hu2016segmentation], aiming to localize video segments corresponding to textual queries. Early approaches fall into two categories: proposal-based methods [gao2017ctrl, anne2017mcn, xu2019qspn, chen2019sap, chen2018tgn, zhang2020tan, liu2024tuning], which rely on pre-generated candidate segments, and proposal-free methods [chen2020gdp, lu2019debug, zeng2020drn, ghosh2019excl, zhang2020vslnet, rodriguez2020proposal, rodriguez2021dori], which directly predict temporal boundaries without proposals. Recent advancements leverage DETR-based architectures [lei2021qvhighlights, moon2023qd-detr, sun2024trdetr, jang2023eatr, lee2025bamdetr, xiao2024uvcom, um2025kwdetr, tang2025simdetr, tang2023cotdet, zhang2025temporally], inspired by their success in object detection, to improve MR performance through end-to-end query modeling. Concurrently, multimodal large language models (MLLMs) [meinardus2024mrblip, lu2024llavamr, huang2024vtimellm, wang2024timerefine, huang2024lita, guo2024vtgllm] exploit cross-modal knowledge to achieve remarkable results by aligning linguistic and visual semantics.
Beyond architectural innovations, data-centric approaches have emerged to address MR challenges. Some works [lu2019debug, chen2020gdp, zeng2020drn, rodriguez2023memory] tackle data imbalance issues or expensive computational costs, while others incorporate large-scale external datasets for pretraining. Notably, VIDEOLIGHTS [paul2024videolights] leverages BLIP-2’s image-to-text generation to synthesize training data. Although effective, these methods often depend on external resources. In contrast, we propose a simpler, generalizable data augmentation strategy that enhances performance without external data or pretrained models.
@@
Video Data Augmentation has been widely explored to address the challenges of limited and difficult-to-annotate video data. Common augmentation techniques involve spatial transformations, such as random flipping and cropping, or temporal modifications, such as random downsampling. In recent years, novel augmentation strategies have emerged. For instance, [yun2020videomix] introduced a method for video classification that randomly mixes videos along spatial and temporal dimensions, assigning ground truth labels proportionally based on voxel distribution. Building on this, [zhang2020selfpaced] employed GAN [goodfellow2020gan] to generate more realistic video data, while [gowda2022learn2augment] proposed a learning-based approach to select optimal video pairs for higher-quality augmentation. Additionally, [pan2024labelefficient] utilized video diffusion models to generate labeled videos for the guidewire segmentation task.
@@ 3 Method
Existing moment retrieval frameworks (e.g., M-DETR [lei2021qvhighlights]) face significant challenges under data scarcity and distribution imbalance. While DETR-based architectures [sun2024trdetr, yang2024taskweave, xiao2024uvcom, lee2025bamdetr] achieve end-to-end localization through encoder-decoder paradigms, their performance is constrained by two critical limitations: 1) High annotation costs for video data result in limited training samples, causing models to overfit specific training distributions; 2) Practical scenarios contain inherently ambiguous event boundaries with adjacent segment overlaps and semantically similar events that induce localization confusion.
@@
Our work focuses on leveraging limited and imbalanced training samples to enhance temporal localization models’ generalization capabilities in real-world scenarios. Consequently, the proposed AMR adheres to the standard vanilla DETR-based framework, comprising feature extraction, transformer encoder, and transformer decoder modules. For feature extraction, we align with previous work by using CLIP [radford2021clip] and SlowFast [feichtenhofer2019slowfast] to extract video features while employing the CLIP text encoder for text feature extraction, where the encoded video features are denoted as $V\in\mathbb{R}^{L\times C}$ ($L$ video frames) and the text features as $T\in\mathbb{R}^{K\times C}$ ($K$ text tokens), with $C$ representing the hidden dimension. In the transformer encoder, cross-attention from visual to text features encodes multimodal video representation $\hat{V}\in\mathbb{R}^{L\times C}$ and text representation $\hat{T}\in\mathbb{R}^{K\times C}$. The decoder processes $N$ learnable queries $Q\in\mathbb{R}^{N\times C}$ with video features through several attention blocks and feed-forward networks, generating updated queries $\hat{Q}\in\mathbb{R}^{N\times C}$. Final boundary coordinates and confidence scores are predicted through bounding box and classification heads.
@@ 3.2 Two-stage Training: Cold Start and Distillation
In traditional single-stage training frameworks [sun2024trdetr, yang2024taskweave, xiao2024uvcom, lee2025bamdetr], the distribution discrepancy between augmented data $\mathcal{D}_{\text{aug}}$ and real data $\mathcal{D}_{\text{real}}$ often leads to a decline in model generalization. Specifically, the sharp boundaries of events in synthetic data may cause the model to over-rely on abrupt visual features (e.g., scene cuts or sudden object appearances) while neglecting gradual transition patterns in real-world scenarios (e.g., natural continuity of human actions). To address this issue, we propose a cold-start distillation two-stage training framework (Fig. 3), whose core idea is to establish fundamental boundary localization and semantic modeling capabilities through augmented data in the cold-start stage and achieve migration to real data via a knowledge distillation strategy in the distillation adaptation stage.
@@ 3.3 Training Objective
Hungarian Matching To establish one-to-one correspondence between predicted and ground-truth objects without manual heuristics, we employ the Hungarian matching algorithm following the DETR framework [carion2020detr]. The matching task is formulated as a minimum-cost bipartite graph matching problem. Specifically, a bipartite graph is constructed between the prediction set (containing $N$ elements) and the padded ground-truth set (expanded to $N$ elements with empty sets). The optimal assignment is determined by minimizing the matching cost:
@@
Loss Function During the cold-start stage training, we utilize the baseline loss functions from M-DETR [lei2021qvhighlights], including the classification loss $\mathcal{L}_{\text{cls}}$, the localization loss $\mathcal{L}_{\text{loc}}$, and the auxiliary saliency loss $\mathcal{L}_{\text{sal}}$ (see [lei2021qvhighlights] for implementation details). The composite loss function is defined as:
@@
M-DETR [lei2021qvhighlights] 53.94 34.84 - - 32.20 52.89 33.02 54.82 29.40 30.73
UMT [liu2022umt] 60.26 44.26 - - 38.59 56.23 41.18 53.83 37.01 36.12
QD-DETR [moon2023qd-detr] 62.68 46.66 62.23 41.82 41.22 62.40 44.98 62.52 39.88 39.86
UniVTG [lin2023univtg] 59.74 - - - 36.13 58.86 40.86 57.60 35.59 35.47
EaTR [jang2023eatr] 61.36 45.79 61.86 41.91 41.74 - - - - -
MomentDiff [li2023momentdiff] - - - - - 57.42 39.66 54.02 35.73 35.95
TR-DETR [sun2024trdetr] 67.10 51.48 66.27 46.42 45.09 64.66 48.96 63.98 43.73 42.62
TaskWeave [yang2024taskweave] 64.26 50.06 65.39 46.47 45.38 - - - - -
UVCOM [xiao2024uvcom] 65.10 51.81 - - 45.79 63.55 47.47 63.37 42.67 43.18
BAM-DETR [lee2025bamdetr] 65.10 51.61 65.41 48.56 47.61 62.71 48.64 64.57 46.33 45.36
@@ 4.1 Datasets and Metrics
Datasets Our study evaluates performance on three benchmark datasets for moment retrieval: QVHighlights [lei2021qvhighlights], Charades-STA [gao2017ctrl], and TACoS [regneri2013tacos]. QVHighlights consists of 10,148 videos, primarily sourced from YouTube, including vlogs and news clips. Notably, this dataset allows a single textual query to correspond to multiple segments within a video, making it more representative of real-world applications. Therefore, we adopt it as our primary benchmark. Charades-STA comprises 9,848 videos, focusing on human activities in indoor scenes. TACoS consists of 127 cooking-related videos, characterized by long video durations and short target segments, posing a significant challenge for moment retrieval.
@@
2D-TAN [zhang2020tan] 58.76 46.02 27.50 41.25 40.01 27.99 12.92 27.22
M-DETR [lei2021qvhighlights] 65.83 52.07 30.59 45.54 37.97 24.67 11.97 25.49
MomentDiff [li2023momentdiff] - 55.57 32.42 - 44.78 33.68 - -
QD-DETR [moon2023qd-detr] - 57.31 32.55 - - - - -
UniVTG [lin2023univtg] 70.81 58.01 35.65 50.10 51.44 34.97 17.35 33.60
UVCOM [xiao2024uvcom] - 59.25 36.64 - - 36.39 23.32 -
@@
Metrics We employ evaluation metrics consistent with prior works [yang2024taskweave, lei2021detecting]. For QVHighlights, we report Recall@1 (R1) at Intersection over Union (IoU) thresholds of 0.5 and 0.7. Given that this dataset allows multiple matching segments per query, we also use mean average precision (mAP) as an evaluation metric, reporting scores at IoU thresholds of 0.5, 0.75, and the average over all IoUs. For Charades-STA and TACoS, we evaluate Recall@1 (R1) at IoU thresholds of 0.3, 0.5, and 0.7, along with the mean IoU (mIoU) of the top-1 predictions.